Find E_y the y-component of the total electric fie
Answer #1

(1) From Pythagorean theorem, Distance between 5nC and 4nC charge is

r_{1}= \sqrt{x_{1}^{2} + y_{3}^{2}}= \sqrt{(-1)^{2}+3^{2}} = 2m

(2) From Pythagorean theorem, Distance between 7nC and 4nC charge is

r_{2}= \sqrt{x_{2}^{2} + y_{3}^{2}}= \sqrt{2^{2}+3^{2}} = \sqrt{13}m

Thanks for using HomeworkLib.com

Do vist again :)